# Flextron: Many-in-One Flexible Large Language Model

### v1

@@ 4.1 Experimental Settings
We perform our evaluation on the GPT3 and Llama2 (Touvron et al., 2023) model families. GPT3 is a representative multilingual model family (Shoeybi et al., 2019) with $2$ and $8$ billion parameter variants (among others); these are pretrained with the NeMo framework (Kuchaiev et al., 2019). The total number of trainable parameters for GPT3-$2$B is $2$ billion, with $1.2$ billion non-embedding parameters. The model contains $24$ layers, with a hidden dimension of $2048$. Each MHA layer possesses $16$ heads. GPT3-$8$B comprises $8$ billion parameters, of which $6.4$ billion are non-embedding parameters. The model contains $32$ layers, each with a hidden dimension of $4096$. Each MHA layer possesses $32$ heads. Both models are trained on $1.1$ trillion tokens, where data is obtained from publicly available data sources, comprising $53$ languages and code. Both models support a maximum context length of $4096$. We further validate our approach using the Llama2-7B model (Touvron et al., 2023), a widely used open-source pre-trained model with $6.5$ billion non-embedding parameters. This model employs a $32$-layer transformer architecture with a hidden dimension of $4096$, incorporating $32$ attention heads for each Multi-Head Attention (MHA) mechanism. We perform neuron/head sorting, elastic continued-training and automatic network selection with the same domain data, unless otherwise specified.
@@ 4.2 Results
Full 6.5 B 75.1% 71.5% 77.5% 69.1% 43.0% 78.1% 69.1%
Static-0.7$\times$ 4.2 B 65.8% 64.2% 75.6% 62.3% 32.5% 67.1% 61.3%
Dynamic-0.7$\times$ 4.1 B 68.6% 65.1% 76.1% 63.7% 35.9% 69.4% 63.1%
Static-0.6$\times$ 4.0 B 66.1% 63.8% 75.0% 62.1% 31.9% 68.0% 61.2%
Dynamic-0.6$\times$ 3.9 B 67.1% 63.8% 74.9% 62.2% 34.5% 69.7% 62.0%
Static-0.5$\times$ 3.5 B 65.9% 61.7% 74.8% 61.9% 30.3% 67.6% 60.4%
Flextron-Llama2-7B Dynamic-0.5$\times$ 3.4 B 66.5% 62.9% 74.1% 62.0% 33.4% 68.5% 61.2%
@@
Table 4: Flextron training costs compared to pretraining cost. We report the number of tokens to illustrate the training cost.
@@ Appendix A Architecture Visualization
We provide searched architectures in Figure 14, based on two variants of GPT3 family. The observation validates our previous heuristic entailed in Sec 5.1.

### v2

@@ 4.1 Experimental Settings
We perform our evaluation on the GPT3 and Llama2 (Touvron et al., 2023) model families. GPT3 is a representative multilingual model family (Shoeybi et al., 2019) with $2$ and $8$ billion parameter variants (among others); these are pretrained with the NeMo framework (Kuchaiev et al., 2019). The total number of trainable parameters for GPT3-$2$B is $2$ billion, with $1.2$ billion non-embedding parameters. The model contains $24$ layers, with a hidden dimension of $2048$. Each MHA layer possesses $16$ heads. GPT3-$8$B comprises $8$ billion parameters, of which $6.4$ billion are non-embedding parameters. The model contains $32$ layers, each with a hidden dimension of $4096$. Each MHA layer possesses $32$ heads. Both models support a maximum context length of $4096$. GPT3 2B and 8B are trained on $1.1$ trillion tokens, where data is obtained from publicly available data sources, comprising $53$ languages and code (Shoeybi et al., 2019). We further validate our approach using the Llama2-7B model (Touvron et al., 2023), a widely used open-source pre-trained model with $6.5$ billion non-embedding parameters. This model employs a $32$-layer transformer architecture with a hidden dimension of $4096$, incorporating $32$ attention heads for each Multi-Head Attention (MHA) mechanism. Llama2-7B is trained on $2$ trillion tokens (Touvron et al., 2023). We perform neuron/head sorting, elastic continued-training and router tuning with the same domain data. For both GPT3 and Llama2, we use $89.9$B tokens for elastic continued training, and $1.049$B tokens for router tuning. For Llama2, we additionally sample a subset comprising of $2.62$B tokens from the Nemotron-4 curated continued training dataset (Parmar et al., 2024) for joint tuning.
@@ 4.2 Results
Full 6.5 B 75.1% 71.5% 77.5% 69.1% 45.1% 78.1% 69.4%
Static-0.7$\times$ 4.2 B 65.8% 64.2% 75.6% 62.3% 41.9% 67.1% 62.8%
Dynamic-0.7$\times$ 4.1 B 68.6% 65.1% 76.1% 63.7% 42.2% 69.4% 64.2%
Static-0.6$\times$ 4.0 B 66.1% 63.8% 75.0% 62.1% 37.7% 68.0% 62.1%
Dynamic-0.6$\times$ 3.9 B 67.1% 63.8% 74.9% 62.2% 39.4% 69.7% 62.8%
Static-0.5$\times$ 3.5 B 65.9% 61.7% 74.8% 61.9% 35.9% 67.6% 61.3%
Flextron-Llama2-7B Dynamic-0.5$\times$ 3.4 B 66.5% 62.9% 74.1% 62.0% 36.8% 68.5% 61.8%
@@
Table 4: Flextron training costs compared to pretraining cost. We report the number of tokens for elastic CT, router tuning and joint tuning (in case of Llama2-7B) to illustrate the training cost.
@@ Appendix A Architecture Visualization
We provide searched architectures in Figure 14, based on two variants of the GPT3 family. The observation validates our previous heuristic entailed in Sec 5.1.